# ViDDAR: Vision Language Model-Based Task-Detrimental Content Detection for Augmented Reality

### v1

@@ 2.1 Obstruction Attack
We begin by modeling obstruction attacks, as they are relatively simpler and more objective to evaluate. This process aligns with the principles of full-reference image quality assessment methods[47], where comparisons are made between the raw image $I_{r}$ and the altered – in our case, augmented, – image $I_{a}$. While users can only see $I_{a}$, the system simultaneously monitors both $I_{r}$ and $I_{a}$. In $I_{r}$, there exists a set $K$ containing $n$key objects," which are potentially important and may need users’ attention. The key objects $k_{i}\in K$ are not predefined and can vary dynamically depending on different scenarios. For instance, while a stop sign is often considered important, it may not be treated as a key object when it appears as a product to be sold on a shelf, as Fig. 3 shows. For each key object $k_{i}$, there is a pixel-level mask $m_{k}^{i}$ which defines the extent of $k_{i}$ as a set of pixels; $|m_{k}^{i}|$ denotes the number of pixels within $m_{k}^{i}$. All pixel-level masks $m_{k}^{i}$ collectively form a set, denoted as $M_{k}$.
@@ 4.1.1 Experiment Setup
Detection Method Vision Language Model Key Object Recognition Accuracy (%) Segmentation Mean IoU (%) Obstruction Attack Detection Accuracy (%)
@@
Figure 8: Statistical analysis of datasets we collected: (a): Distribution of the key object’s obstructed area percentage for "obstructed" and "not obstructed" images in the obstruction dataset; (b): Label distribution percentages for attributes $A$ (alignment precision), $S$ (style similarity), $I$ (information misrepresentation), and $M$ (information manipulation) in the information manipulation dataset.
@@ 4.2.1 Experiment Setup
Figure 9: Information manipulation attack dataset samples. The first row shows raw images; the second row shows the augmented images. Their labels ($Alignment\ Precision$, $Style\ Similarity$, $Information\ Misrepresentation$, $Information\ Manipulation$) for each sample are as follows: (a): A virtual plant on the speaker may lead to information manipulation, making the speaker appear to be a plant pot, with labels (1, 1, 1, 1); (b): The virtual exit sign placed above a door is potentially misleading, but it is not well aligned with the door, labeled as (0, 1, 1, 0); (c) A virtual coffee cup on a laptop may give the impression that the laptop is a food tray, potentially causing damage if other food is placed on it. However, the low-quality texture of the coffee cup makes it more noticeable, resulting in labels (1, 0, 1, 0); (d) The toy dinosaur is well aligned with the ground and the trash bin and has a high-quality texture, but it is unlikely to misrepresent the function of the trash bin, thus labeled (1, 1, 0, 0).
@@
{mdframed}
"Here are two images. The first one is a raw image, and the second one is an augmented image, created by adding some virtual content to the space.
Please answer the following questions:
1. What is the virtual content in the augmented image?
2. What ‘key object’ is interacting with the virtual content? Avoid general terms like ‘table surface’ or ‘environment.’
3. Is the virtual content accurately aligned to the object, without a significant gap? Answer yes or no, then explain why.
4. Does the virtual content have a relatively high-quality yet reasonable texture that blends it into the real world? Answer yes or no, then explain why.
5. Do you think the interaction will make users believe the ‘key object’ has some false functionality or information it does not have, or lose some true functionality or information it actually has? Think creatively, only say no if the combination has no specific semantic relation. Answer yes or no, then explain why.
6. If you answered ‘yes’ in all questions 3, 4, and 5, you must say ‘yes.’ Otherwise, you say ‘no.’"
@@ 4.2.2 Results
Vision Language Model Detection Accuracy Detection Precision Detection Recall
@@ 4.3.2 Results
Figure 10: An example of statements used in the information manipulation dataset validation study. The four statements correspond to information misrepresentation ($I$), style similarity ($S$), alignment precision ($A$), and overall manipulation labeling ($M$), respectively.
@@ 5 Limitations and Future Work
Detection Task Network Settings Mean Latency
Obstruction Detection One-hop 533 ms
Six-hop 960 ms
Information Manipulation Detection One-hop 9.62 s
Six-hop 12.30 s

### v2

@@ 2.1 Obstruction Attack
We begin by modeling obstruction attacks, as they are relatively simpler and more objective to evaluate. This process aligns with the principles of full-reference image quality assessment methods[47], where comparisons are made between the raw image $I_{r}$ and the altered – in our case, augmented, – image $I_{a}$. While users can only see $I_{a}$, the system simultaneously monitors both $I_{r}$ and $I_{a}$. In $I_{r}$, there exists a set $K$ containing $n$ "key objects," which are potentially important and may need users’ attention. The key objects $k_{i}\in K$ are not predefined and can vary dynamically depending on different scenarios. For instance, while a stop sign is often considered important, it may not be treated as a key object when it appears as a product to be sold on a shelf, as Fig. 3 shows. For each key object $k_{i}$, there is a pixel-level mask $m_{k}^{i}$ which defines the extent of $k_{i}$ as a set of pixels; $|m_{k}^{i}|$ denotes the number of pixels within $m_{k}^{i}$. All pixel-level masks $m_{k}^{i}$ collectively form a set, denoted as $M_{k}$.
@@ 4.1.1 Experiment Setup
Detection Method
Vision Language
Model
Key Object Recognition
Accuracy (%)
Segmentation
Mean IoU (%)
Obstruction Attack
Detection Accuracy (%)
@@
Figure 8: Statistical analysis of datasets we collected: (a): Distribution of the key object’s obstructed area percentage for "obstructed" and "not obstructed" images in the obstruction dataset; (b): Label distribution percentages for attributes AA (alignment precision), $S$ (style similarity), II (information misrepresentation), and $M$ (information manipulation) in the information manipulation dataset.
@@ 4.2.1 Experiment Setup
Figure 9: Information manipulation attack dataset samples. The first row shows raw images; the second row shows the augmented images. Their labels (AlignmentPrecisionAlignment\ Precision, StyleSimilarityStyle\ Similarity, InformationMisrepresentationInformation\ Misrepresentation, InformationManipulationInformation\ Manipulation) for each sample are as follows: (a): A virtual plant on the speaker may lead to information manipulation, making the speaker appear to be a plant pot, with labels (1, 1, 1, 1); (b): The virtual exit sign placed above a door is potentially misleading, but it is not well aligned with the door, labeled as (0, 1, 1, 0); (c) A virtual coffee cup on a laptop may give the impression that the laptop is a food tray, potentially causing damage if other food is placed on it. However, the low-quality texture of the coffee cup makes it more noticeable, resulting in labels (1, 0, 1, 0); (d) The toy dinosaur is well aligned with the ground and the trash bin and has a high-quality texture, but it is unlikely to misrepresent the function of the trash bin, thus labeled (1, 1, 0, 0).
@@
"Here are two images. The first one is a raw image, and the second one is an augmented image, created by adding some virtual content to the space. Please answer the following questions: 1. What is the virtual content in the augmented image? 2. What ‘key object’ is interacting with the virtual content? Avoid general terms like ‘table surface’ or ‘environment.’ 3. Is the virtual content accurately aligned to the object, without a significant gap? Answer yes or no, then explain why. 4. Does the virtual content have a relatively high-quality yet reasonable texture that blends it into the real world? Answer yes or no, then explain why. 5. Do you think the interaction will make users believe the ‘key object’ has some false functionality or information it does not have, or lose some true functionality or information it actually has? Think creatively, only say no if the combination has no specific semantic relation. Answer yes or no, then explain why. 6. If you answered ‘yes’ in all questions 3, 4, and 5, you must say ‘yes.’ Otherwise, you say ‘no.’"
@@ 4.2.2 Results
Vision Language
Model
Detection
Accuracy
Detection
Precision
Detection
Recall
@@ 4.3.2 Results
Figure 10: An example of statements used in the information manipulation dataset validation study. The four statements correspond to information misrepresentation ($I$), style similarity ($S$), alignment precision (AA), and overall manipulation labeling ($M$), respectively.
@@ 5 Limitations and Future Work
Detection Task
Network Settings
Mean Latency
Obstruction
Detection
One-hop
533 ms
Six-hop
960 ms
Information
Manipulation
Detection
One-hop
9.62 s
Six-hop
12.30 s